A conducting wire has a 1.2 mm diameter, a 2.8 m length, and a 70 m? resistance. What is the resistivity of the material?

I will give 5 stars for the correct answer
Answer #1

R = rho*L/A

==> rho = R*A/L

= 70*10^-3*(pi*1.2^2*10^-6/4)/2.8

= 2.83*10^-8 ohm-m
Answer #2

Resistance of a wire in ?
R = ?L/A
? is resistivity of the material
L is length in meters
A is cross-sectional area in m
Answer #3

Here

r = 1.2/2 = 0.6 mm

L = 2.8 m

R = 70 mohm

R = 0.070 Ohm

Now ,


.070 = p*2.8/(3.141*.0006^2)

p = .070*(3.141*.0006^2)/2.8

p = 2.83 *10^-8 Ohm . m

the resistivity of the material is 2.83 *10^-8 Ohm . m